OpenStudy (anonymous):

Pretty please help! >< Evaluate the expression when r = 36. 272 + r – 78 ???

pooja195 (pooja195):

Hi welcome to openstudy :-) Plug in the value of r 272+36-78=? Baby steps ;p 272+36=?

OpenStudy (anonymous):

272+36=308 and lol

pooja195 (pooja195):

Yay.